# Teaching Large Language Models to Reason through Learning and Forgetting

### v2

@@ 5.2 Boosting Chain-of-Thought Reasoning
We present CoT reasoning results at inference time for LLMs fine-tuned on the Countdown (2(b)) and the Game-of-24 datasets (Table 3). Similar to Ye et al. (2024), we also evaluate the models trained on Countdown on the Game-of-24 test set, because conceptually Game-of-24 is a subset of Countdown. For the Tables 2(b) and 3, we use the highest learning rate (1e-5 for Q1.5B and 5e-6 for Q7B) that yields best performance in CoT reasoning, leaving further discussion on the learning rate for Sec. 5.3. For the UFT objective, we fix the coefficient for the same model across different datasets, selected by the average success rate on validation set .
As noted in Sec. 4.2, prior work on SFT using synthetic data (Yuan et al., 2023; Singh et al., 2023; Yang et al., 2024) typically consider learning correct data generated by CoT only. Therefore, in both the Tables 2(b) and 3, only the gray cells (SFT on CoT data) represent the baselines, and all the other cells involve novelty of our approach (search data sources and/or UFT). We analyze these in detail below.
Training data quality is decisive to CoT reasoning.
@@
First, we find that the test success rates are mainly determined by the training data quality, outweighing the effect of training objective (SFT or UFT). Data quality is measured by the best-of-n success rate: fraction of unique solved cases out of all the training cases.777Although ToT and RAP greatly outperform CoT as inference-time methods (Table 1, Table 5) in Game-of-24, they produce lower-quality training data than CoT (Table 3), because CoT’s best-of-n (n=100) success rate far exceeds its average (93% versus 6% in Q7B, 82% versus 5% in Q1.5B). The best performances in Table 3 and 2(b) correspond to the datasets with the highest quality (CoT data in Game-of-24, BFS+DFS data in Countdown). In addition, when the two datasets have similar quality (CoT vs CoT+ToT+RAP in Game-of-24, BFS+DFS vs CoT+BFS+DFS in Countdown), their performances are also alike. Remarkably, fine-tuning on the high-quality classic search (BFS+DFS) data from Countdown yields significantly better performance on Game-of-24 than models fine-tuned directly on Game-of-24 CoT data (37% vs 28.2% for Q7B and 29.7% vs 14.5% for Q1.5B). This highlights the importance of incorporating high-quality training data, rather than relying solely on CoT data, as is common practice.
Unlikelihood auxiliary loss helps CoT reasoning.
Next, we analyze the effect of incorporating UL loss (Eq. 5) to fine-tuning. By comparing SFT and UFT cells, we find UL helps improve performance in most scenarios: 14 out of 18 times for Q1.5B and 16 out of 18 for Q7B. While UFT provides a smaller boost than improving training data quality (typically under 4%), it still offers notable gains. For instance, UFT increases the success rate of Q7B fine-tuned on all data from 27.5% to 30.2%, surpassing its best inference-time baseline (28%). It also yields a substantial 7% improvement on the Game-of-24 test set for Q7B trained on all Countdown data.
@@ 5.3 Mitigating Forgetting in Search-based Reasoning
We investigate whether fine-tuned LLMs can retain search-based reasoning abilities (e.g., ToT, RAP), a question that remains underexplored in the absence of continued pretraining. Evaluating this retention is analogous to measuring backward transfer (Lopez-Paz & Ranzato, 2017) in continual learning, i.e., whether performance on a prior task (search-based reasoning) degrades after fine-tuning on a new one (CoT-style data). We focus on Q7B, which exhibits strong search-based reasoning performance and a clear gap over CoT baselines (Table 1). For completeness, the results of Q1.5B is shown in Appendix C.
Although learning rate decay is employed following the alignment handbook, we find that using a peak learning rate of 2e-5, as in the original Qwen2.5-Math setup (Yang et al., 2024), leads to catastrophic forgetting (McCloskey & Cohen, 1989), with near-zero performance in inference-time search. We hypothesize that this degradation stems from the lack of reward signals in CoT-style supervision, which weakens the model’s reward modeling ability essential for search-based reasoning (see Sec. A.4 for details).
@@
Inspired by continual learning literature (Mirzadeh et al., 2020), we adopt a much smaller learning rate to prevent parameters from drifting too far from their initial state. Fig. 3 and Fig. 5 shows the effects of (peak) learning rate in standard SFT on Q7B. As we reduce the learning rate from 5e-6 to 1e-6 (an exceptionally low value for SFT), we observe a clear trade-off: CoT reasoning performance declines rapidly (though it remains above the baseline), while inference-time search improves significantly (though often remains below the baseline). To our knowledge, this is a novel finding in the LLM literature, where learning rate is rarely considered as a simple yet effective way to mitigate forgetting.
@@
Building on the findings in Fig. 3, we observe that the effect of a low learning rate extends to UFT as well. Here, we examine the impact of the unlikelihood loss specifically at the lowest learning rate of 1e-6. Consistent with Sec. 5.2, UFT provides marginal improvements over SFT in CoT reasoning, regardless of the choice of . However, with =1e-4, inference-time search improves considerably compared to SFT (=0) in most cases. For example, with the same CoT data, ToT achieves 28.6% vs. 22.2% under SFT, and RAP reaches 28.9% vs. 17.5% – far exceeding the RAP baseline of 10.2%. This indicates that learning to forget incorrect CoT-style paths could generalize to forgetting incorrect intermediate search steps.
@@ 6 Conclusion
We propose a simple and effective method for mathematical reasoning, addressing three underexplored challenges: (1) unifying mixed-format data from diverse reasoners into a CoT-style format, (2) leveraging unpaired correct and failed paths via unlikelihood fine-tuning, and (3) mitigating loss of search capability using a small learning rate. Experiments on Countdown and Game-of-24 show that data quality, unlikelihood loss, and learning rate are key to balancing CoT efficiency and search capability.
Appendix A Reasoner Details
A.1 The Problems
@@
We randomly split the first 900 cases (rank #1-900 by human average performance in the dataset888https://github.com/maitrix-org/llm-reasoners/blob/main/examples/ToT/game24/data/24.csv) into (720 cases) and (180 cases). The are the next 100 cases (rank #901-1000) following the setup in ToT (Yao et al., 2023a). As a result, the validation set evaluates in-distribution (ID) generalization, while the test set evaluates out-of-distribution (OOD) generalization based on human difficulty levels. Each case is guaranteed to be solvable using the four input numbers, which range from 1 to 13.
@@
We follow SoS codebase999https://github.com/kanishkg/stream-of-search (Gandhi et al., 2024) to randomly generate 500k training cases, 1k validation cases, and 1k test cases. The training and validation sets share the same distribution of target numbers, while the test set includes distinct target numbers (including the number 24). Each problem consists of four input numbers sampled from 1 to 99, with a target number between 10 and 100. All generated problems are guaranteed to have at least one valid solution path using only integer intermediate results.
A.2 CoT Reasoner
@@ A.3 ToT and RAP Reasoners
A.4 Loss of Search Capability after CoT Fine-Tuning: A Planning Perspective
@@ A.5 Classic BFS and DFS Reasoners
Appendix B Fine-tuning Details
We follow the alignment handbook101010https://github.com/huggingface/alignment-handbook (Tunstall et al., 2024) to fine-tune two base LLMs, Qwen2.5-Math 1.5B (Q1.5B)111111https://huggingface.co/Qwen/Qwen2.5-Math-1.5B and Qwen2.5-Math 7B (Q7B)121212https://huggingface.co/Qwen/Qwen2.5-Math-7B (Yang et al., 2024).
We extend the training codebase to support Unlikelihood Fine-Tuning (UFT), where UFT and Supervised Fine-Tuning (SFT) differ only by the unlikelihood loss coefficient . A batch size is used for both successful and failed data batches (after gradient accumulation). We disable sequence packing to prevent cross-contamination131313https://github.com/huggingface/transformers/issues/25452 and set a context length of 256 tokens. which is sufficient for our tasks. Training is performed in bfloat16 precision, and model checkpoints are saved every 5% epochs for later evaluation.
The learning rate follows cosine_with_min_lr schedule141414https://github.com/huggingface/transformers/blob/a22a4378d97d06b7a1d9abad6e0086d30fdea199/src/transformers/optimization.py#L338C5-L338C48: it linearly warms up from 0 to over the first 10% epochs, then decays to a minimum of 7e-8 via cosine function for the rest epochs. In Sec. 5.2, we use a peak learning rate of 1e-5 for Q1.5B and 5e-6 for Q7B, while in Sec. 5.3, we decrease to much lower values, as discussed in the main paper.
The training time is measured in number of epochs based on the correct dataset size . Although training runs for multiple epochs, each failed path is shown to the model only once, as the pool of failed examples is significantly larger than the correct dataset.
For Countdown, BFS+DFS generate 614k correct paths. The full CoT+BFS+DFS dataset contains 734k correct paths for Q1.5B and 985k for Q7B. Due to the large dataset size, we train for 2 epochs. The training time on CoT+BFS+DFS data is less than 5 hrs for Q1.5B and 30 hrs for Q7B on an instance with 8 A100s (40 GB). For Game-of-24, CoT+ToT+RAP contains 6k correct paths for Q1.5B and 14k for Q7B. We train for 10 epochs, with total training time under 0.5 hrs for Q1.5B and 2.5 hrs for Q7B.
Appendix C Additional Results
Limited effects on 1.5B due to weak base performance.
Although not the primary focus of this work, we provide results for Q1.5B in Appendix (Fig. 6 to Fig. 8) for completeness. Given its weak search-based reasoning performance as a base model (Table 5; success rate 2% in Countdown), the effects of learning rate, CoT data, and unlikelihood loss are minimal, likely due to its poor initial capability. This suggests that the observed benefits may be more pronounced at larger model scales.

### v3

@@ 5.2 Boosting Chain-of-Thought Reasoning
We present CoT reasoning results at inference time for LLMs fine-tuned on the Countdown (2(b)) and the Game-of-24 datasets (3(b)). Similar to Ye et al. (2024), we also evaluate the models trained on Countdown on the Game-of-24 test set, because conceptually Game-of-24 is a subset of Countdown. For both tables, we use the highest learning rate (1e-5 for Q1.5B and 5e-6 for Q7B) that yields best performance in CoT reasoning, leaving further discussion on the learning rate for Sec. 5.3. For the UFT objective, we fix the coefficient for the same model across different datasets, selected by the average success rate on validation set . Our key findings are summarized below.
NLL loss is essential for offline fine-tuning on reasoning tasks.
The most salient observation is that SimPO, while successful in alignment tasks using only a preference loss (Meng et al., 2024b), yields near-zero success across model sizes and datasets. In contrast, CPO-SimPO (Xu, 2024), which simply adds an NLL term to SimPO, achieves substantially better results. This highlights the critical role of NLL loss, shared by SFT, CPO-SimPO, and UFT, which explicitly encourages models to follow correct reasoning paths. These findings align with prior results in Pang et al. (2024); Pal et al. (2024).
Incoporating high-quality training data is decisive for CoT reasoning.
@@
Second, among methods that use NLL loss, performance is primarily determined by the quality of the training data, outweighing the effect of forgetting objective. We measure data quality by the best-of-n success rate: the fraction of unique solved cases out of all the training cases.777Although ToT and RAP greatly outperform CoT as inference-time methods (Table 1, Table 6) in Game-of-24, they produce lower-quality training data than CoT (3(b)), because CoT’s best-of-n (n=100) success rate far exceeds its average (93% versus 6% in Q7B, 82% versus 5% in Q1.5B). The best performances in 2(b) and 3(b) correspond to the datasets with the highest quality (BFS+DFS data in Countdown, CoT data in Game-of-24). In addition, when the two datasets have similar quality (BFS+DFS vs CoT+BFS+DFS in Countdown, CoT vs CoT+ToT+RAP in Game-of-24), their performances are also alike. Remarkably, fine-tuning on the high-quality classic search (BFS+DFS) data from Countdown yields significantly better performance on Game-of-24 than models fine-tuned directly on Game-of-24 CoT data (37% vs 28.2% for Q7B and 29.7% vs 14.5% for Q1.5B). This highlights the importance of incorporating high-quality training data, rather than relying solely on CoT data, as is common practice.
Unpaired learning and forgetting is more robust than paired one.
Lastly, we analyze the effect of forgetting loss to SFT. First, we observe that UFT consistently improves over SFT in most scenarios: 14 out of 18 times for Q1.5B and 16 out of 18 for Q7B. In terms of magnitude, UFT yields an average improvement of 1.3% (with a maximum of 4%) in 3(b), and an average of 1.5% (with a maximum of 7%) in 2(b). In contrast, CPO-SimPO shows mixed results. While it improves over SFT with an average gain of 1.2% and maximum of 5% in 3(b), it consistently underperforms SFT in all scenarios in 2(b), with an average drop of 3.5%.
We believe several factors contribute to the contrasting behavior of CPO-SimPO relative to SFT (and UFT) when fine-tuned on Countdown versus Game-of-24. First, as shown in Appendix Table 5, the paired preference requirement in CPO-SimPO results in the exclusion of 0.1% to 5.8% of correct data on Countdown, while no such filtering occurs for Game-of-24. Since correct data determines data quality, this reduction likely contributes to CPO-SimPO’s weaker performance on Countdown. Second, the exclusion of unpaired failed data, along with the challenge of balancing preference and NLL objectives, may have a greater impact on the more difficult Countdown task.
@@ 5.3 Mitigating Forgetting in Search-based Reasoning
We investigate whether fine-tuned LLMs can retain search-based reasoning abilities (e.g., ToT, RAP), a question that remains underexplored in the absence of continued pretraining. Evaluating this retention is analogous to measuring backward transfer (Lopez-Paz & Ranzato, 2017) in continual learning, i.e., whether performance on a prior task (search-based reasoning) degrades after fine-tuning on a new one (CoT-style data). We focus on Q7B, which exhibits strong search-based reasoning performance and a clear gap over CoT baselines (Table 1). For completeness, the results of Q1.5B is shown in Appendix B.
Although learning rate decay is employed following the alignment handbook, we find that using a peak learning rate of 2e-5, as in the original Qwen2.5-Math setup (Yang et al., 2024), leads to catastrophic forgetting (McCloskey & Cohen, 1989), with near-zero performance in inference-time search. We hypothesize that this degradation stems from the lack of reward modeling in CoT-style supervision, which would be essential for search-based reasoning (see Sec. C.4 for details).
@@
Inspired by continual learning (Mirzadeh et al., 2020), we adopt a much smaller learning rate to keep parameters within the “basic capability basin” (Chen et al., 2025), preserving pretrained knowledge. Fig. 3 and Fig. 5 shows the effects of (peak) learning rate in standard SFT on Q7B. As we reduce the learning rate from 5e-6 to 1e-6 (an exceptionally low value for SFT), we observe a clear trade-off: CoT reasoning performance declines rapidly (although it remains above the baseline), while inference-time search improves significantly (although often remains below the baseline). This behavior is similar to learning rate effects seen in reference-free preference optimization (Meng et al., 2024b).
@@
Building on the findings in Fig. 3, we observe that the effect of a low learning rate extends to UFT and CPO-SimPO as well. Here, we examine the impact of the unlikelihood loss specifically at the lowest learning rate of 1e-6. Consistent with Sec. 5.2, UFT provides marginal improvements over SFT in CoT reasoning, regardless of the choice of . However, with =1e-4, inference-time search improves considerably compared to SFT (=0) in most cases. For example, with the same CoT data, ToT achieves 28.6% vs. 22.2% under SFT, and RAP reaches 28.9% vs. 17.5% – far exceeding the RAP baseline of 10.2%. This indicates that learning to forget incorrect CoT-style paths could generalize to forgetting incorrect intermediate search steps.
@@ 6 Conclusion
We propose a simple and effective method for mathematical reasoning, addressing three underexplored challenges: (1) augmenting training data by unifying mixed data formats from diverse reasoners into a CoT-style format, (2) leveraging unpaired correct and failed paths via unlikelihood fine-tuning, and (3) mitigating loss of search capability using a small learning rate. Experiments on Countdown and Game-of-24 show that data quality, unlikelihood loss, and learning rate are key to balancing CoT efficiency and search capability.
Appendix A Fine-tuning Details
We follow the alignment handbook888https://github.com/huggingface/alignment-handbook (Tunstall et al., 2024) to fine-tune two base LLMs with full parameters, Qwen2.5-Math 1.5B (Q1.5B)999https://huggingface.co/Qwen/Qwen2.5-Math-1.5B and Qwen2.5-Math 7B (Q7B)101010https://huggingface.co/Qwen/Qwen2.5-Math-7B (Yang et al., 2024). Below are common setups for all the methods (i.e., SFT, UFT, and preference-based methods).
We disable sequence packing to prevent cross-contamination111111https://github.com/huggingface/transformers/issues/25452 and set a context length of 256 tokens. which is sufficient for our tasks. Training is performed in bfloat16 precision, and model checkpoints are saved every 5% epochs for later evaluation.
The learning rate follows cosine_with_min_lr schedule121212https://github.com/huggingface/transformers/blob/a22a4378d97d06b7a1d9abad6e0086d30fdea199/src/transformers/optimization.py#L338C5-L338C48: it linearly warms up from 0 to over the first 10% epochs, then decays to a minimum of 7e-8 via cosine function for the rest epochs. In Sec. 5.2, we use a peak learning rate of 1e-5 for Q1.5B and 5e-6 for Q7B, while in Sec. 5.3, we decrease to much lower values, as discussed in the main paper.
A.1 SFT Baseline and UFT
Following Sec. 4.2, in our implementation, Unlikelihood Fine-Tuning (UFT) and Supervised Fine-Tuning (SFT) differ only by the unlikelihood loss coefficient . A batch size is used for both successful and failed data batches (after gradient accumulation) in UFT.
A.2 Preference-based (Reference-free) Baselines: SimPO and CPO-SimPO
Background.
In addition to SFT, we compare UFT against offline preference-based methods: SimPO (Meng et al., 2024b) and CPO-SimPO (Xu, 2024), both designed to align LLMs with human preferences. While UFT is not a preference-based method per se, our reasoning task can be interpreted through the lens of preference alignment, where successful reasoning paths are preferred over failed ones. Like preference-based approaches, UFT learns from both successful and failed paths, making these methods natural baselines for comparison. The implementations of these methods follow the TRL library131313https://huggingface.co/docs/trl/en/cpo_trainer (von Werra et al., 2020).
SimPO and CPO-SimPO build on DPO (Rafailov et al., 2023) and CPO (Xu et al., 2024), requiring preference pairs but eliminating the need for a reference model used in KL regularization (i.e., reference-free). This results in improved training efficiency and reduced GPU memory usage. By contrast, while KTO (Ethayarajh et al., 2024) like UFT operates on unpaired data, it still requires a reference model during training. In our preliminary experiments, running KTO using the same global batch size as UFT necessitated precomputing reference model logits and offloading the optimizer to the CPU, resulting in a increase in training time on the Game-of-24 dataset and out-of-memory errors on the Countdown dataset. Consequently, we exclude KTO from our baselines and instead focus on reference-free approaches such as SimPO.
Table 4 summarizes the comparison between UFT and several offline preference-based methods. Notably, UFT is both reference-free and does not require preference pairs.
Offline fine-tuning method | Preference data | Reference-free? |
DPO (Rafailov et al., 2023), (non-iterative) RPO (Pang et al., 2024) | Paired | No |
CPO (Xu et al., 2024), SimPO (Meng et al., 2024b), CPO-SimPO (Xu, 2024) | Paired | Yes |
KTO (Ethayarajh et al., 2024) | Unpaired | No |
UFT (Ours) | Unpaired | Yes |
Simple Preference Optimization with a Reference-Free Reward (SimPO) (Meng et al., 2024b).
SimPO relies on paired preference data, denoted as , where are successful and failed reasoning paths for the same input , respectively. To construct from our original dataset , we first filter for inputs that have both successful and failed paths. For each successful path , we generate preference pairs by randomly sampling failed paths corresponding to the same input, treating them as rejected responses. This sampling strategy ensures that the resulting dataset is similar in size to that used for UFT and captures as much diversity as possible within the constraint of pairing.
The training objective is defined as:
where is the sigmoid function, is a scaling factor, and is the target reward margin, following default values in TRL. The length normalization terms (, ) help reduce bias toward longer responses. This normalization could also be explored in UFT as future work. The margin enforces a minimum separation between the preferred and rejected log-probabilities.
CPO-SimPO (Xu, 2024).
Although pure preference-based methods (e.g., DPO, SimPO) perform well in alignment tasks, they have been found to underperform on reasoning tasks that require explicit maximization of preferred responses (Meng et al., 2024b; Pal et al., 2024; Pang et al., 2024). While CPO-SimPO was not originally motivated by reasoning tasks, we adopt it because CPO-SimPO addresses this issue by augmenting SimPO with the NLL term from CPO (Xu et al., 2024):
where the NLL coefficient following the TRL.
Connection with RPO (Pang et al., 2024).
CPO-SimPO (Xu, 2024) resembles a non-iterative version of Iterative Reasoning Preference Optimization (IRPO) (Pang et al., 2024), which we call RPO:
where is the reference model (i.e., the base LLM). CPO-SimPO becomes equivalent to RPO when a reference model is introduced, the length normalization and reward margin in the preference term are removed, and length normalization is applied in the NLL term instead. However, since RPO, like DPO and KTO, depends on a reference model, we exclude it from our baseline comparisons.
Appendix B Additional Results
B.1 Training Dataset Statistics
Training data (task, base, reasoner) | Total correct paths | Paired correct paths | Quality |
Countdown, Q7B, CoT | 428.0k | 97.50% | 37.6% |
Countdown, N/A, BFS+DFS | 614.2k | 97.55% | 86.2% |
Countdown, Q7B, CoT+BFS+DFS | 985.3k | 99.88% | 86.9% |
Countdown, Q1.5B, CoT | 147.9k | 94.22% | 20.0% |
Countdown, Q1.5B, CoT+BFS+DFS | 734.3k | 99.86% | 86.3% |
Game-of-24, Q7B, CoT | 9.1k | 100% | 92.9% |
Game-of-24, Q7B, ToT+RAP | 5.1k | 100% | 69.3% |
Game-of-24, Q7B, CoT+ToT+RAP | 13.7k | 100% | 95.3% |
Game-of-24, Q1.5B, CoT | 4.4k | 100% | 82.1% |
Game-of-24, Q1.5B, ToT+RAP | 1.7k | 100% | 44.7% |
Game-of-24, Q1.5B, CoT+ToT+RAP | 6.0k | 100% | 87.6% |
For Countdown, we train for 2 epochs due to the large dataset size, shown as in Table 5. On an instance with 8×A100 GPUs (40GB), total training time on CoT+BFS+DFS data is under 5 hours for Q1.5B and under 30 hours for Q7B using UFT. For Game-of-24, we train for 10 epochs. Total training time is under 0.5 hours for Q1.5B and under 2.5 hours for Q7B using UFT. SimPO and CPO-SimPO only require a single dataloader over paired data, resulting in roughly 2× shorter training time compared to UFT.
B.2 Results on Qwen2.5-Math 7B
B.3 Results on Qwen2.5-Math 1.5B
Although not the primary focus of this work, we provide fine-tuning results based on Qwen2.5-Math 1.5B in Fig. 6 and Fig. 7 for completeness. Given its weak search-based reasoning performance as a base model (e.g., success rate 2% in Countdown) shown in Table 6, the effects of learning rate, CoT data, and unlikelihood loss are minimal, likely due to its poor initial capability. This suggests that the observed benefits depend on the model size.
Test set & Inference | Base LLM |
Countdown (1000 cases) | |
CoT (greedy) | 2.2% / 0.6m |
CoT (pass@1) | 2.1% / 3.2m |
search: ToT | 2.2% / 68m |
search: RAP | 1.4% / 143m |
Game-of-24 (100 cases) | |
CoT (greedy) | 2.0% / 0.1m |
CoT (pass@1) | 3.0% / 0.5m |
search: ToT | 8.0% / 5.4m |
search: RAP | 4.0% / 11.3m |
Appendix C Reasoner Details
C.1 The Problems
@@
We randomly split the first 900 cases (rank #1-900 by human average performance in the dataset141414https://github.com/maitrix-org/llm-reasoners/blob/main/examples/ToT/game24/data/24.csv) into (720 cases) and (180 cases). The are the next 100 cases (rank #901-1000) following the setup in ToT (Yao et al., 2023a). As a result, the validation set evaluates in-distribution (ID) generalization, while the test set evaluates out-of-distribution (OOD) generalization based on human difficulty levels. Each case is guaranteed to be solvable using the four input numbers, which range from 1 to 13.
@@
We follow SoS codebase151515https://github.com/kanishkg/stream-of-search (Gandhi et al., 2024) to randomly generate 500k training cases, 1k validation cases, and 1k test cases. The training and validation sets share the same distribution of target numbers, while the test set includes distinct target numbers (including the number 24). Each problem consists of four input numbers sampled from 1 to 99, with a target number between 10 and 100. All generated problems are guaranteed to have at least one valid solution path using only integer intermediate results.
C.2 CoT Reasoner
@@ C.3 ToT and RAP Reasoners
C.4 Loss of Search Capability after CoT Fine-Tuning: A Planning Perspective